# Direct Adaptive Control of Grid-Connected Power Converters via Output-Feedback Data-Enabled Policy Optimization

### v1

@@ III-B DeePO for direct adaptive LQR control
$\nabla J_{t}(V)=2\left(\overline{U}_{0,t}^{\top}R\overline{U}_{0,t}+\overline{X}_{1,t}^{\top}P_{t}\overline{X}_{1,t}\right)V\Sigma_{t}(V),$ (9)
@@
$P_{t}=Q+V^{\top}\overline{U}_{0,t}^{\top}R\overline{U}_{0,t}V+V^{\top}\overline{X}_{1,t}^{\top}P_{t}\overline{X}_{1,t}V.$
DeePO is direct and adaptive in the sense that it directly uses online closed-loop data to update the policy. Thus, it can rapidly adapt to dynamic changes reflected in the data. Algorithm 1 has a recursive policy update and can be implemented efficiently. Specifically, all covariance matrices and the inverse $\Phi_{t+1}^{-1}$ have recursive updates. Moreover, the parameterization can be updated recursively via rank-one update, i.e.,
@@
where $\phi_{t}=[u_{t}^{\top},x_{t}^{\top}]^{\top}$, and $\Phi_{t}^{-1}$ and $V_{t}^{\prime}$ are given from the last iteration. Theoretically, it is shown that under mild assumptions the policy $\{K_{t}\}$ converges to the optimal LQR gain. We refer to [24, Section IV] for detailed discussions.
The DeePO framework developed so far considers the state feedback setting. However, the state of power converter system (1) is not measurable, and only input-output data is available for control design. Such an system is typically referred to as partially observed [29]. Next, we extend DeePO to partially observed systems for adaptive control using input-output data.
Algorithm 1 DeePO for direct adaptive LQR control
1:Offline data $(X_{0,t_{0}},U_{0,t_{0}},X_{1,t_{0}})$, an initial policy $K_{t_{0}}$, and a stepsize $\eta$.
2:for $t=t_{0},t_{0}+1,\dots$ do
3:     Apply $u_{t}=K_{t}x_{t}+e_{t}$ and observe $x_{t+1}$.
4:     Update covariance matrices $\Phi_{t+1}$ and $\overline{X}_{1,t+1}$.
5:     Policy parameterization: given $K_{t}$, solve $V_{t+1}$ via
@@
6:     Update of the parameterized policy: perform one-step projected gradient descent
$V_{t+1}^{\prime}=V_{t+1}-\eta\Pi_{\overline{X}_{0,t+1}}\nabla J_{t+1}(V_{t+1}),$ (10)
where the gradient $\nabla J_{t+1}(V_{t+1})$ is given by Lemma 1.
7:     Gain update: update the control gain by
@@
8:end for
IV Output-feedback DeePO for partially observed systems
In this section, we extend DeePO for direct adaptive control of partially observed systems using input-output data.
Consider the system (1). Since the state is unmeasurable, we represent (1) with a non-minimal realization using input and output signals. Denote the input trajectory, the output trajectory, and their stack from time $t-n$ to $t-1$ as
${u}_{t,n}=\begin{bmatrix}u_{t-1}\\ \vdots\\ u_{t-n}\end{bmatrix},~{}{y}_{t,n}=\begin{bmatrix}y_{t-1}\\ \vdots\\ y_{t-n}\end{bmatrix},~{}{z}_{t}=\begin{bmatrix}{u}_{t,n}\\ {y}_{t,n}\end{bmatrix}$
respectively. Define the observability matrix
$\mathcal{O}=\begin{bmatrix}CA^{n-1}\\ \vdots\\ CA\\ C\end{bmatrix},$
the controllability matrices ${}\mathcal{C}=[B~{}~{}AB~{}~{}\cdots~{}~{}A^{n-1}B]$, $\mathcal{C}_{w}=[I~{}~{}A~{}~{}\cdots~{}~{}A^{n-1}],$ and the Toeplitz matrices capturing the impulse and disturbance response
$\mathcal{T}=\begin{bmatrix}0&CB&CAB&\cdots&CA^{n-2}B\\ 0&0&CB&\cdots&CA^{n-3}B\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ 0&\cdots&&0&CB\\ 0&0&0&0&0\end{bmatrix},$
$\mathcal{T}_{w}=\begin{bmatrix}0&C&CA&\cdots&CA^{n-2}\\ 0&0&C&\cdots&CA^{n-3}\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ 0&\cdots&&0&C\\ 0&0&0&0&0\end{bmatrix}.$
Let $S=[C(\mathcal{C}-A^{n}O^{\dagger}\mathcal{T}),CA^{n}O^{\dagger}]$, $d_{t}=C(\mathcal{C}_{w}-A^{n}\mathcal{O}^{\dagger}\mathcal{T}_{w})w_{t,n}-CA^{n}\mathcal{O}^{\dagger}v_{t,n}+v_{t}$. Then, we have the following results, the proof of which is provided in the appendix.
Theorem 1 (Non-minimal controllable realization)
A non-minimal controllable realization of (1) is given by (LABEL:equ:nonnimimal_realization) shown at the bottom of the next page.
Remark 1
The length of the past trajectory used to represent the state does not necessarily match the system order $n$. In fact, it only needs to exceed the system lag (i.e., the observability index), under which the dimension of $z_{t}$ can be reduced.
Then, we can apply DeePO for the system (LABEL:equ:nonnimimal_realization), whose state is measurable. Define the matrix of input-output data
$Z_{0,t}=\begin{bmatrix}z_{0}&z_{1}&\dots&z_{t-1}\end{bmatrix}\in\mathbb{R}^{n(m+d)\times t}.$
Assume that the data is PE, i.e., the block matrix $[U_{0,t}^{\top},Z_{0,t}^{\top}]^{\top}$ has full row rank. Then, the covariance parameterization of $K$ for the system (LABEL:equ:nonnimimal_realization) is given by
$\begin{bmatrix}K\\ I_{n(m+d)}\end{bmatrix}=\frac{1}{t}\begin{bmatrix}U_{0,t}\\ Z_{0,t}\end{bmatrix}\begin{bmatrix}U_{0,t}\\ Z_{0,t}\end{bmatrix}^{\top}V=:\Psi_{t}V,$
where $V\in\mathbb{R}^{n(2m+d)\times n(m+d)}$.
Define the sample covariance matrices
$\overline{Z}_{0,t}=\frac{1}{t}Z_{0,t}\begin{bmatrix}U_{0,t}\\ Z_{0,t}\end{bmatrix}^{\top},~{}\overline{Z}_{1,t}=\frac{1}{t}Z_{1,t}\begin{bmatrix}U_{0,t}\\ Z_{0,t}\end{bmatrix}^{\top}.$
Denote the realization (LABEL:equ:nonnimimal_realization) as $(A_{z},B_{z},C_{z})$. Following the certainty-equivalence principle and disregarding $d_{t}$, we approximate the closed-loop matrix by
$A_{z}+B_{z}K\approx\overline{Z}_{1,t}V.$
Then, the LQR problem of (LABEL:equ:nonnimimal_realization) is parameterized as
$\mathop{\text{minimize}}\limits_{V}~{}J_{t}(V):=\text{Tr}\left((Q_{1}+V^{\top}\overline{U}_{0,t}^{\top}R_{z}\overline{U}_{0,t}V)\Sigma_{t}(V)\right),$ (11)
subject $\text{subject to}~{}~{}\overline{Z}_{0,t}V=I_{n},$
where $Q_{z},R_{z}$ are positive definite matrices with proper dimensions, $\Sigma_{t}(V)=I_{n}+\overline{Z}_{1,t}V\Sigma_{t}(V)V^{\top}\overline{Z}_{1,t}^{\top}$, and the gain matrix can be recovered by $K=\overline{U}_{0,t}V$.
The DeePO method for partially observed systems is provided in Algorithm 2. Next, we apply Algorithm 2 to stabilize power converter and renewable energy systems.
Algorithm 2 DeePO for partially observed systems
1:Offline data $(Z_{0,t_{0}},U_{0,t_{0}},Z_{1,t_{0}})$, an initial policy $K_{t_{0}}$, and a stepsize $\eta$.
2:for $t=t_{0},t_{0}+1,\dots$ do
3:     Apply $u_{t}=K_{t}z_{t}+e_{t}$ and observe $z_{t+1}$.
4:     Update covariance matrices $\Phi_{t+1}$ and $\overline{Z}_{1,t+1}$.
5:     Policy parameterization: given $K_{t}$, solve $V_{t+1}$ via
$V_{t+1}=\Psi_{t+1}^{-1}\begin{bmatrix}K_{t}\\ I_{n}\end{bmatrix}.$
6:     Update of the parameterized policy: perform one-step projected gradient descent
$V_{t+1}^{\prime}=V_{t+1}-\eta\Pi_{\overline{Z}_{0,t+1}}\nabla J_{t+1}(V_{t+1}).$
7:     Gain update: update the control gain by
$K_{t+1}=\overline{U}_{0,t+1}V_{t+1}^{\prime}.$
8:end for
@@ V DeePO for stabilization of power converters and renewable energy systems
In this section, we first discuss the implementation of DeePO on the power converter systems. Then, we perform simulations for a grid-connected power converter and a direct-drive wind generator.
V-A Implementation of DeePO for power converter systems
We consider two typical power converter systems: a grid-connected power converter (with lithium batteries as the DC source) in Fig. 1 and a direct-drive (type 4) wind generator in Fig. 3, which also uses a converter as the grid interface. Both systems share similar control challenges, including unknown models, measurement limitations, and potential change of dynamics. While Algorithm 2 considers only time-invariant systems, it can potentially handle changes in the system dynamics. This is because DeePO uses online closed-loop data to update the policy in real time and can quickly adapt to changes. Moreover, since Algorithm 2 has a recursive implementation, the online update of the policy is computationally efficient.
For both systems, we select the deviations of active and reactive power from their reference values as the outputs of the unknown system, denoted by $y=[P_{E}-P_{0},Q_{E}-Q_{0}]^{\top}$. The DeePO method is configured to provide two control inputs that are added to the current references of the system. All input and output signals are expressed in per-unit ($\mathrm{p.u.}$) values. Since directly measuring internal state variables is challenging, we construct new state variables using input-output information as $z_{t}=[u_{t,T_{\text{ini}}}^{\top},y_{t,T_{\text{ini}}}^{\top}]^{\top}$, where the length of the initial trajectory for constructing new states is set to $T_{\text{ini}}=6$ for both systems.
@@ V-A Implementation of DeePO for power converter systems
  • sampling frequency: The sampling frequency is chosen as $200\,\mathrm{Hz}$.
  • trajectory length $T$: The trajectory length is set to $300$ for the power converter and $600$ for the wind generator to satisfy the persistency of excitation condition.
  • weighting matrices: The input weighting matrix is set as $R=I_{m}$, while the output weighting matrix is set as $Q=\text{diag}(I_{mT_{\text{ini}}},100I_{pT_{\text{ini}}})$.
  • gradient descent parameters: The stepsize for gradient descent is set as $\eta=0.005$. For the grid-connected power converter, the number of gradient descent steps per control sampling time is set to $1$. For the wind generator, two configurations are compared with $1$ and $10$ gradient descent steps per control sampling time.
@@ V-B Simulations on a grid-connected power converter
After the oscillation is observed, we inject band-limited white noise signals into the system through the two input channels from $t=2.0\,\mathrm{s}$ to $t=3.5\,\mathrm{s}$ to excite the system and collect data. The DeePO controller is activated at $t=3.5\,\mathrm{s}$, providing real-time control inputs that effectively counteract the oscillation. As depicted in Fig. 2, DeePO effectively damps the oscillation and restores the stable operation of the power converter. In contrast, the responses of the power converter without DeePO are shown as the grey lines in Fig. 2, where the system exhibits sustained active and reactive power oscillations, indicating a low stability margin and increasing the risk of equipment damage and potential cascading failures in the power system.
@@ V-C Simulations on a direct-drive wind generator
Figure 4: The time-domain responses of the wind generator: (a) active power and (b) reactive power. The DeePO is activated at $t=6.0\,\mathrm{s}$. with DeePO (1 gradient descent step per control cycle); with DeePO (10 gradient descent steps per control cycle); without DeePO.
@@
Then, we inject band-limited white noise to collect input-output trajectory data from $t=3.0\,\mathrm{s}$ to $t=6.0\,\mathrm{s}$. At $t=6.0\,\mathrm{s}$, DeePO is activated and successfully provides real-time control inputs, eliminating the oscillations in both configurations, as shown in Fig. 4. In the configuration with fewer gradient descent steps per control cycle (blue line), the convergence of active and reactive power is slower, with acceptable stabilization performance. By comparison, using $10$ steps (orange line) enables faster damping of oscillations. This illustrates a trade-off between computational efficiency and control performance: increasing the gradient steps improves convergence but requires higher computational effort. By comparison, the wind generator system exhibits a low stability margin without DeePO (grey line in Fig. 4), with persistent oscillations in active and reactive power after $t=6.0\,\mathrm{s}$. These persistent oscillations endanger the wind generator itself (as mechanical resonance may be triggered) and also the secure operation of power systems. We observe that the damping performance of DeePO is similar to the performance of DeePC (c.f. Fig. 14 in [21]), but the computational burden is significantly less (since DeePC requires solving a quadratic program online).
@@
In time-varying environments, non-adaptive controllers that rely solely on historical data may fail to maintain effective control. In contrast, DeePO adapts by continuously learning from real-time data, allowing it to better accommodate time-varying system conditions. Fig. 5 shows the time-domain responses of the wind generator using a non-adaptive controller (a fixed controller that achieves similar performance to DeePO before $t=9.0\,\mathrm{s}$) and the DeePO controller, respectively, under parameter changes and disturbances. Before $t=9.0\,\mathrm{s}$, both controllers have converged to their respective optimal gains based on previously collected data. At $t=9.0\,\mathrm{s}$, the DC voltage control parameters are intentionally reduced to 0.9 times of their original values, simulating a sudden change in the wind generator system. While this change does not affect the equilibrium point, it modifies the dynamic characteristics of the system. From $t=9.0\,\mathrm{s}$ to $t=21.0\,\mathrm{s}$, the DeePO controller continues to update its control strategy by leveraging real-time data, whereas the non-adaptive controller retains its original control settings without adaptation. To ensure that the collected data are informative to update the DeePO controller, noise injection is increased between $t=10.0\,\mathrm{s}$ and $t=17.0\,\mathrm{s}$ to better excite the system and capture its dynamic characteristics.
At $t=21.0\,\mathrm{s}$, a small disturbance is introduced by injecting a q-axis voltage perturbation into the PLL. As shown in Fig. 5, the DeePO controller quickly damps out oscillations, demonstrating significantly better response performance compared to the non-adaptive controller. This improvement can be attributed to the real-time adaptation of the control policy in DeePO, which enables it to perceive changes in system dynamics and maintain superior performance under disturbances.
@@ VI Conclusion
In this paper, we proposed a data-enabled policy optimization (DeePO) method, which is direct data-driven, adaptive, based on online input-output data, and recursive in terms of implementation. We demonstrated via simulations the effectiveness of DeePO in stabilizing and mitigating undesired oscillations in a grid-connected power converter and a direct-drive wind generator. In particular, we observed that DeePO has satisfactory performance in stabilization, and shows a similar damping ratio to data-enabled predictive control (DeePC). Nevertheless, DeePO is computationally more efficient than DeePC, as it is gradient-based and does not require solving optimization problems online.
Future work includes the extension of DeePO to achieve other control objectives (e.g., reference tracking) and to other types of systems (e.g., slowly time-varying systems).
Appendix A Proof of Theorem 1
The state can be represented with system dynamics and history trajectories as
$x_{t}$ $=A^{n}x_{t-n}+\mathcal{C}{u}_{t,n}+\mathcal{C}_{w}w_{t,n}$ (12)
${y}_{t,n}$ $=\mathcal{O}x_{t-n}+\mathcal{T}{u}_{t,n}+\mathcal{T}_{w}w_{t,n}+v_{t,n}.$
Since $\mathcal{O}$ has full column rank, it has a unique left pseudo inverse $\mathcal{O}^{\dagger}=(\mathcal{O}^{\top}\mathcal{O})^{-1}\mathcal{O}^{\top}$. Then, it follows immediately from (12) that
$x_{t}$ $=(\mathcal{C}-A^{n}\mathcal{O}^{\dagger}\mathcal{T}){u}_{t,n}+A^{n}\mathcal{O}^{\dagger}{y}_{t,n}$
$~{}~{}~{}+(\mathcal{C}_{w}-A^{n}\mathcal{O}^{\dagger}\mathcal{T}_{w})w_{t,n}-A^{n}\mathcal{O}^{\dagger}v_{t,n},$
and $y_{t}=Cx_{t}+v_{t}=Sz_{t}+d_{t}.$ Thus, a non-minimal realization of (1) is given by (LABEL:equ:nonnimimal_realization).
The proof for controllablity of (LABEL:equ:nonnimimal_realization) follows the same vein of that of the key reachability lemma in [30, Lemma 3.4.7]. Note that their relatively prime condition holds for the system (1) since the length of the past trajectory used to represent the state matches the system order $n$.

### v2

@@ IV-B Output-feedback DeePO
$\nabla J_{t}(V)=2\left(\overline{U}_{0,t}^{\top}R\overline{U}_{0,t}+\overline{Z}_{1,t}^{\top}P_{t}\overline{Z}_{1,t}\right)V\Sigma_{t},$ (15)
@@
$P_{t}=Q+V^{\top}\overline{U}_{0,t}^{\top}R\overline{U}_{0,t}V+V^{\top}\overline{Z}_{1,t}^{\top}P_{t}\overline{Z}_{1,t}V.$
By Lemma 3, the gradient can be computed by solving two Lyapunov equations. The stepsize $\eta_{t}$ should be chosen based on the signal-to-noise ratio (SNR) of the online data. Specifically, when the SNR is high, the gradient direction is more reliable, allowing for a larger stepsize. Conversely, when the SNR is low, a smaller stepsize is necessary to prevent the policy from deviating from the stability region. Accordingly, we set the stepsize as
$\eta_{t}=\frac{\eta_{0}}{\left\|\overline{U}_{0,t}\Pi_{\overline{Z}_{0,t}}\overline{U}_{0,t}^{\top}\right\|},~{}t\geq t_{0},$ (16)
where $\eta_{0}$ is a constant, and the denominator is used to quantify the SNR.
The DeePO algorithm is direct and adaptive in the sense that it directly uses online closed-loop data to update the policy. Thus, it can rapidly adapt to dynamic changes reflected in the data. Algorithm 1 has a recursive policy update and can be implemented efficiently. Specifically, all covariance matrices and the inverse $\Phi_{t+1}^{-1}$ have recursive updates. Moreover, the parameterization can be updated recursively via rank-one update, i.e.,
@@
where $\phi_{t}=[u_{t}^{\top},z_{t}^{\top}]^{\top}$, and $\Phi_{t}^{-1}$ and $V_{t}^{\prime}$ are given from the last iteration. Theoretically, it is shown that under mild assumptions the policy $\{K_{t}\}$ converges to the optimal LQR gain. We refer to [24, Section IV] for detailed discussions.
Algorithm 1 Output-feedback DeePO
1:Offline data $(u_{-\bar{l}},y_{-\bar{l}},\dots,u_{t_{0}-1},y_{t_{0}-1},y_{t_{0}})$ and a stepsize $\eta_{t}$.
2:Computation of the permutation matrix: Constitute the data matrix $\Xi_{0,t_{0}}$ in (7) and perform SVD
$\Xi_{0,t_{0}}=\begin{bmatrix}U_{r}&U_{l}\end{bmatrix}\begin{bmatrix}\Lambda_{r}&0\\ 0&\Lambda_{l}\end{bmatrix}\begin{bmatrix}V_{r}^{\top}\\ V_{l}^{\top}\end{bmatrix},$
where $r>m\bar{l}$ is selected such that there is a clear distinction between the largest $r$ singular values and the remaining ones. Let $T=\Lambda_{r}^{-1}U_{r}^{\top}.$
3:Compute the initial policy $K_{t_{0}}$ from (14) with offline data.
4:for $t=t_{0},t_{0}+1,\dots$ do
5:     Compute $z_{t}=T\xi_{t}$, apply $u_{t}=K_{t}z_{t}+e_{t}$ to the system, and observe $z_{t+1}$.
6:     Policy parameterization: given $K_{t}$, solve $V_{t+1}$ via
@@
7:     Update of the parameterized policy: perform one-step projected gradient descent
$V_{t+1}^{\prime}=V_{t+1}-\eta_{t}\Pi_{\overline{Z}_{0,t+1}}\nabla J_{t+1}(V_{t+1}).$
8:     Gain update: update the control gain by
@@
9:end for
@@ V DeePO for stabilization of power converters and renewable energy systems
In this section, we first discuss the implementation of output-feedback DeePO on the power converter systems. Then, we perform simulations for a grid-connected power converter and a direct-drive wind generator.
V-A Implementation of output-feedback DeePO for power converter systems
We consider two typical power converter systems: a grid-connected power converter (with lithium batteries as the DC source) in Fig. 1 and a direct-drive (type 4) wind generator in Fig. 3, which also uses a converter as the grid interface. The maximum lags for the two systems are $\bar{l}=2$ and $4$, respectively. Both systems share similar control challenges, including unknown models, measurement limitations, and potential change of dynamics. While Algorithm 1 considers only time-invariant systems, it can potentially handle changes in the system dynamics. This is because DeePO uses online closed-loop data to update the policy in real time and can quickly adapt to changes. Moreover, since Algorithm 1 has a recursive implementation, the online update of the policy is computationally efficient.
For both systems, the outputs of the unknown dynamics are defined as the deviations of active and reactive power from their reference values: $y=[P_{E}-P_{0},Q_{E}-Q_{0}]^{\top}$. The DeePO method is configured to provide two control inputs added to the current references of the system. All input and output signals are expressed in per-unit ($\mathrm{p.u.}$) values. Since internal states are not directly measurable, the input-output stack is constructed as $\xi_{t}=[u_{t,\bar{l}}^{\top},y_{t,\bar{l}}^{\top}]^{\top}$. The matrix $\Xi_{0,t}$ is built from past trajectories, followed by SVD to determine the reduced dimension $r$. A projection matrix $T$ is then computed to obtain the reduced-order state $z_{t}$ from $\xi_{t}$.
@@ V-A Implementation of output-feedback DeePO for power converter systems
  • The sampling frequency is chosen as $200\,\mathrm{Hz}$.
  • The trajectory length is set to $300$ for the power converter and $600$ for the wind generator.
  • The input and state weighting matrices are set as $R=I_{m}$ and $Q=I_{r}$, respectively.
  • The stepsize for gradient descent is set according to (16) with $\eta_{0}=0.0001$.
@@ V-B Simulations on a grid-connected power converter
After the oscillation is observed, we inject band-limited white noise signals into the system through the two input channels from $t=2.0\,\mathrm{s}$ to $t=3.5\,\mathrm{s}$ to excite the system and collect data. We then construct the Hankel matrix $\Xi_{0,t}$ using the input-output data from this interval and perform SVD. The resulting singular values, in descending order, are 7.075, 2.596, 0.556, 0.496, 0.476, 0.460, 0.249, 0.1867.0752.5960.5560.4960.4760.4600.2490.1867.075,\ 2.596,\ 0.556,\ 0.496,\ 0.476,\ 0.460,\ 0.249,\ 0.1867.075 , 2.596 , 0.556 , 0.496 , 0.476 , 0.460 , 0.249 , 0.186. The minimum embedding dimension is $m\bar{l}=4$. Since the first six singular values exhibit a clear distinction from the remaining two, we select $r=6$. The projection matrix $T$ is then computed based on the leading components, and the reduced-order state is obtained accordingly. This representation is used to solve an SDP problem of (14), yielding $K_{t_{0}}$ as the initial policy for DeePO. The DeePO controller is activated at $t=3.5\,\mathrm{s}$, providing real-time control inputs that effectively counteract the oscillation. As depicted in Fig. 2, DeePO effectively damps the oscillation and restores the stable operation of the power converter. In contrast, the responses of the power converter without DeePO are shown as the grey lines in Fig. 2, where the system exhibits sustained oscillations, indicating a low stability margin and increasing the risk of equipment damage and potential cascading failures in the power system.
@@ V-C Simulations on a direct-drive wind generator
Figure 4: The time-domain responses of the wind generator: (a) active power and (b) reactive power. The DeePO is activated at $t=6.0\,\mathrm{s}$. with DeePO; without DeePO.
@@
Then, we inject band-limited white noise to collect input-output trajectory data from $t=3.0\,\mathrm{s}$ to $t=6.0\,\mathrm{s}$. Following the same procedure as in the last subsection, we determine the embedding dimension as $r=12$ for this case. Based on the data collected in this interval, we solve an SDP of (14) to obtain the initial DeePO policy $K_{t_{0}}$. As shown in Fig. 4, DeePO is activated at $t=6.0\,\mathrm{s}$ and successfully provides real-time control inputs, eliminating the oscillations. By comparison, the wind generator system exhibits a low stability margin without DeePO, with persistent oscillations in active and reactive power after $t=6.0\,\mathrm{s}$. These sustained oscillations not only threaten the integrity of the wind generator itself, due to potential mechanical resonance, but also compromise the secure operation of the overall power system. We observe that the damping performance of DeePO is comparable to that of DeePC (c.f. Fig. 14 in [21]). Nevertheless, DeePO offers a significant computational advantage. For instance, on an AMD Ryzen 7 PRO 7840H CPU with 32 GB of RAM, solving a quadratic program online (required by DeePC) takes approximately $15~{}\mathrm{ms}$ per time step, whereas DeePO only requires around $1~{}\mathrm{ms}$ to perform a gradient step. This highlights the efficiency of DeePO in delivering similar control performance at a much lower computational cost.
@@
In time-varying environments, non-adaptive controllers that rely solely on historical data may fail to maintain effective control. To illustrate this, we compare a non-adaptive controller with DeePO under changing system conditions. Both controllers apply the same initial policy $K_{t_{0}}$ up to $t=9.0\,\mathrm{s}$. At this point, the DC voltage control parameters are reduced to 90% of their nominal values, altering the system dynamics without affecting the equilibrium point. After $t=9.0\,\mathrm{s}$, the non-adaptive controller retains the fixed gain $K_{t_{0}}$, while DeePO updates its policy online using real-time input-output data. Fig. 5 shows the time-domain responses under this time-varying condition and a subsequent disturbance introduced at $t=21.0\,\mathrm{s}$ by injecting a q-axis voltage perturbation into the PLL. While the non-adaptive controller exhibits degraded performance, DeePO successfully maintains stability and quickly damps out oscillations. This improvement stems from its ability to continuously adapt to changing dynamics in real time.
@@ VI Conclusion
In this paper, we proposed a output-feedback data-enabled policy optimization (DeePO) method, which is direct data-driven, adaptive, based on online input-output data, and recursive in terms of implementation. We demonstrated via simulations the effectiveness of DeePO in stabilizing and mitigating undesired oscillations in a grid-connected power converter and a direct-drive wind generator.
Future work includes the extension of the output-feedback DeePO to achieve other control objectives (e.g., reference tracking) and to other types of systems (e.g., slowly time-varying systems). It is also valuable to investigate the effect of regularization on output-feedback DeePO of power converter systems [30].